# Do We Need Subsidiarity in Software?

### v1

@@ Abstract.
subsidiarity, data privacy
submissionid: 3093ccs: Human-centered computing Empirical studies in HCIccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
We apply subsidiarity in a technological context in the following way: control of digital spaces should be made at the lowest level possible and the highest level necessary. This understanding of subsidiarity encompasses all of the kinds of control a user can have over their device: from customizing one’s desktop experience through methods akin to “Linux Ricing” (Cowell, 2016) to content moderation on a decentralized social media platform. While there are many different ways a user can hold or surrender control in the context of today’s digital technologies, we choose to focus on one particular application of subsidiarity in this paper: subsidiarity in terms of data privacy. In this way, we evaluate the tradeoffs between privacy and convenience called into question with surveillance capitalism today.
@@
  • RQ1: To what extent is the software we use every day are sending and receiving data beyond expected or necessary usage? In other words, what level of control is the software currently operating at?
@@
  • The empirical contributions of our data flow analysis, which provide insight into how various applications could be transmitting our personal data
  • The empirical contributions of our user interviews, which provide insight into how computer users view the tradeoffs between personal privacy and convenience
@@
Beyond furthering the conversation regarding the ethics of data privacy, we hope this paper will demonstrate the principle of subsidiarity as an effective critical lens through which to promote user flourishing in technology design. The principle of subsidiarity can help designers concerned with user well-being promote user autonomy as much as possible, while also being able to discern and justify instances of higher level control for the benefit of the user.
@@ 2. Related Work
We situate our research within a number of areas of prior work. First, we look at previous work in human-computer interaction that engages with the principle of subsidiarity or similar frameworks. Next, we consider previous research in data sharing and tracking, especially in the contexts of devices and browsers. Finally, we consider prior research of user perceptions and attitudes towards data privacy. In all cases, we highlight the unique contributions of our work.
@@ 2.1. The Principle of Subsidiarity and Similar Frameworks in HCI and Beyond
Two prior works in HCI have considered the principle of subsidiarity explicitly. Both demonstrated subsidiarity as an effective principle for re-imagining more ethical social media. First, Hasinoff et al. shows subsidiarity can help improve the balance between context and scale in content moderation (Hasinoff and Schneider, 2022). We build on this work by bringing subsidiarity into a new domain and through conducting empirical studies. In their work on common good-oriented design patterns for social media, Conwill et al. argued that designing for the principle of subsidiarity would encourage more respectful and authentic conversations, and help to resist exploitative mechanisms from tech companies (Conwill et al., 2025). They also note that subsidiarity entails allowing users to customize their experiences more, giving users greater control over the content displayed in their feeds and the amount of notifications they receive. In this work we expand the notion of control into the realm of data privacy.
The principle of subsidiarity has also been proposed as a critical lens for LLM development. Specifically, a recent issue of the Red Hat Research Quarterly proposed that subsidiarity should be a guiding principle for open-source AI development. Amid debates of how “open” open-source models should be, the authors propose that subsidiarity can help determine how much access different communities have to model components based on how much control of the model they need for their use case (Craven et al., 2025). In this way, subsidiarity is upheld when control is given to the most localized community that is able to effectively make their own decisions.
Related to subsidiarity in HCI is the Anarchist HCI movement (Keyes et al., 2019). Social anarchists hold that “human dignity is greatest when human autonomy is greatest” and that communities are the most appropriate decision-makers and loci of power (Keyes et al., 2019). In this we see resonance with subsidiarity, which also holds that autonomy supports human dignity and that lower levels of decision-making are more just. Previous contributions within anarchist HCI proposed how anarchism can influence HCI as a discipline, offering ways to re-think the contributions of our works and how HCI researchers relate to their study participants and the broader HCI research community (Keyes et al., 2019). In contrast, our work considers how subsidiarity can influence the design and use of the computer technologies themselves.
@@ 2.2. Data Sharing and Tracking at the Device Level and Browser Level
Prior work in both HCI and other computer science disciplines has investigated privacy and data tracking on different types of devices. These studies have found that various types of devices — fertility trackers (Hudig and Singh, 2025), Apple mobile devices (Bourdoucen and Lindqvist, 2024), and Android devices (Gamba et al., 2020) — access and share data in ways unexpected to the user. This data sharing spans countries and organizations, including to advertising and tracking services (Hudig and Singh, 2025; Gamba et al., 2020). We extend this prior work both by investigating the data flows from different operating systems, specifically, Windows and Ubuntu, and analyzing our findings through the critical lens of the principle of subsidiarity. We draw methodologically from the work of Hudig et al. (Hudig and Singh, 2025), incorporating both data flow monitoring and user interviews in our work.
@@
Google, Facebook, Microsoft, and Twitter are the most dominant third-party trackers, especially in Western countries, with Google being the significantly biggest offender (Schelter and Kunegis, 2016; Gill et al., 2013; Roesner et al., 2012; Lerner et al., 2016; Karaj et al., 2018; Englehardt and Narayanan, 2016; Krishnamurthy and Wills, 2009; Cassel et al., 2022). More recently, TikTok has become a major third-party tracker (Munir et al., 2023). We will rely on these findings to help determine which platforms are “known data harvesters,” which is one piece of our determining the level of control of an application.
Mainstream browsers experience 70% more third-party tracking and advertising requests on average compared to privacy-focused browsers and experience fingerprinting on significantly fewer websites (Cassel et al., 2022). As such, we include one privacy-focused browser, Brave, in our investigation.
We additionally build on this prior research by investigating the endpoints contacted not only in the contexts of the device/operating system level and the browser, but also during the use of various web applications.
@@ 3. Theory: The Principle of Subsidiarity
In this section we will define the principle of subsidiarity, outline its history and use, and argue why we believe it is an effective critical lens through which to envision a better relationship with technology in the future.
@@
The principle of subsidiarity is a principle of social organization that holds that decisions should be made at the most local level possible and the highest level necessary. In other words, local governing bodies should be empowered to make decisions for themselves, and central authorities should only perform tasks that cannot be done at a local level (for Justice and Peace, 2004). The principle of subsidiarity promotes human dignity through a prioritization of human agency and autonomy: by encouraging decisions to be made at the lowest level possible, individuals and communities are able to make decisions about matters that affect them most, fostering a sense of responsibility and participation (Benedict XVI, 2009). However, subsidiarity is not libertarianism: an important element of subsidiarity is acknowledging the importance of higher levels of intervention when necessary (for Justice and Peace, 2004). With that said, according to subsidiarity, higher levels should only intervene with justification of their necessity.
@@
Historically, the principle of subsidiarity is perhaps most prominently associated with the Christian Democratic political ideology, having its roots in both Catholic Social Teaching (for Justice and Peace, 2004) and the neo-Calvinist teaching of sphere sovereignty (sub, [n. d.]). However, its resonance goes beyond Christian Democrats alone. For example, today subsidiarity may be most well-known as a general principle of European Union law (sub, [n. d.]). Importantly, the principle of subsidiarity resonates with both the political left and the political right: this broad resonance makes it worthy of our consideration. The political left finds resonance with the principle of subsidiarity because it upholds participatory government, resists centralized economic and political power, and can support self-determination for historically marginalized groups. Subsidiarity resonates with the political right because it can justify limited government and supports giving primacy to traditional institutions like the family in addressing moral questions.
@@
While typically associated with questions of political organization, the principle of subsidiarity can apply in other contexts. For example, in the workplace, a boss can empower her employees through the principle of subsidiarity by avoiding micro-managing. In this way, the gifts and talents of the employees are honored by giving the employee a greater ability to contribute freely.
@@
The principle of subsidiarity can offer much in the space of ethical technology design as well. Importantly, Catholic Social Teaching — the Catholic Church’s doctrine about social justice and the primary framework through which subsidiarity emerged — began in response to the societal concerns of the Industrial Revolution (for Justice and Peace, 2004). Specifically, subsidiarity emerged as a response to increasing centralization of economic power (for Justice and Peace, 2004), a trend we see repeating today with “big tech.” While systems like surveillance capitalism undermine personal autonomy, subsidiarity aims to reclaim it. Yet, because subsidiarity guides the discernment of when intervention from higher levels of power is necessary, it can help to tease out the tension between user desires for privacy and the convenience of giving up their data, and help to clarify when data sharing is appropriate or necessary. As mentioned in the related work, subsidiarity has been identified in multiple social media studies as a principle that could guide the design of healthier online spaces due to its prioritization of operating on a smaller scale, and it has also been proposed as a guiding principle in open-sourcing large language models. We are excited to bring the critical lens of subsidiarity into a new technical realm in this paper. Related to previous work in anarchist HCI (Keyes et al., 2019), the principle of subsidiarity also plays a key role in the Catholic Worker Movement: an anarchist Catholic social justice movement that lives out Catholic Social Teaching in an embodied way. Specifically, the principle of subsidiarity helps to define the Catholic Worker Movement’s understanding of anarchy (Linder, 2024). In this way our work is rooted in and builds on previous critical lenses employed in HCI.
@@
As mentioned in the introduction, subsidiarity in a technical context means control of digital spaces should be made at the lowest level possible and the highest level necessary. In the specific application to data privacy, subsidiarity means that data collection should occur at the lowest level possible and highest level necessary for effective use of the application. We define the levels of control we consider in this paper from lowest to highest as follows: user, community, and platform. We define control at the user level to mean that data collection should be transparent and require informed consent to be collected and used in ways not immediately obvious from the primary function of the application. Control at the platform level means that companies can collect, share, and use our data how they please, typically in exchange for the conveniences afforded by them collecting and using our data. Finally, control at the community level pertains to communities, employers, organizations, or special interest groups having control of their participant’s data, typically for the greater benefit of that community. This may include an employer providing software to its employees, a high school providing Chromebooks to its students through the Google for Education program, or a community-run online forum for a special interest or hobby, among other examples.
@@ 4.1. Data Flow Monitoring
We monitored the network transmissions of different types of everyday software platforms in order to infer the extent to which these platforms are collecting and sharing data. Although it is impossible to get a full picture of what is going on, our aim was to get as much of a sense as we could of the entities being contacted and the potential endpoints where user data could be being transmitted. To do so, we used the network analysis tool Wireshark (wir, [n. d.]) to capture the network activity while using different software platforms. We investigated the network activity of multiple operating systems and browsers on their own, as well as various web applications, run in a particular browser on a particular operating system.
@@ 4.1.1. Operating System and Browser Selection
We wanted to compare operating systems and browsers that are more conventional — i.e., owned by large tech companies, who do not make many promises about privacy — to operating systems and browsers that are open-source and/or claim to better protect user privacy. Thus, we selected the operating systems Microsoft Windows and Ubuntu Linux, and the browsers Google Chrome, Firefox, and Brave (bra, [n. d.]). Brave has been demonstrated in the literature to be the most effective browser at reducing third-party tracking and advertising requests (Cassel et al., 2022).
@@ 4.1.3. Data Collection
The social networking experiments ran as follows: open the browser, navigate to the site, log in, scroll for one minute, log out, close the browser. The protocol was slightly different for YouTube: for that experiment we did not log in, we searched for “Taylor Swift,” clicked on the first video, and watched for one minute.
Finally, the chat experiments ran as follows: open the browser, navigate to the site, log in, send a few messages, and close the browser.
@@ 4.1.4. Data Analysis
Through data flow monitoring, because we are observing from the outside, we cannot know if data is only sent to the application for the purpose the user intended (e.g., sending email data to a mail application for the purpose of sending an email, no other data is collected from the user, and the data from the email itself is not used for any other purposes), or if the data sent for an intended purpose is additionally used for purposes unintended by the user. Additionally, it is harder to distinguish between community and platform levels of control in data flow monitoring, because that is determined more by who the data is being shared with rather than that data is shared at all. Thus, we cannot definitively determine if the software operates at user, community, or platform levels of control. Instead, we determine if the software operates at a lower (closer to user-level) or higher (closer to platform or community-level) of control through the following criterion: is application data being sent to the parent company (or any other company) while the software platform is running idle? Additionally, are the platforms only contacting IP addresses affiliated with their parent company and/or known cloud service providers, or are they contacting unexpected IPs (e.g., advertisers or trackers)? Finally, is the parent company a known data harvester (e.g., Google, Meta)? If the software platforms we examine are doing any of these things we determine that they are operating at a higher level of control; if none, they operate at a lower level of control. The flow chart in Figure 1 outlines our decision-making path for the level of control a software application operates at.
@@
We determined both the IPs being contacted during a particular experiment and the amount of application data being sent to various IPs through the Wireshark logs. We extracted the IP addresses from the Wireshark logs and used the reverse IP address lookup tool IPAPI (ipa, [n. d.]) to determine the organizations associated with those IP addresses. We focused on the destination IP addresses, that is, the IPs being contacted by the software. We were unable to read the contents of the packets due to encryption, which is a common challenge in work on network monitoring (Hudig and Singh, 2025).
@@
Finally, for the activity experiments, because we would expect application data to be transferred to the parent company while the user is performing an activity, we primarily focus on if the IPs contacted are expected or unexpected. For the unexpected IPs, we additionally investigated how much application data was transferred to these endpoints, and report this in the Appendix.
4.2. Interviews
4.2.1. Participant Recruitment

### v2

@@ Abstract.
subsidiarity, catholic social teaching, data privacy
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791098isbn: 979-8-4007-2278-3/2026/04submissionid: 3093ccs: Human-centered computing Empirical studies in HCIccs: Security and privacy Social aspects of security and privacy
@@ 1. Introduction
We apply subsidiarity in a technological context in the following way: control of digital spaces should be made at the lowest level possible and the highest level necessary. This understanding of subsidiarity encompasses all of the kinds of control a user can have over their device: from customizing one’s desktop experience through methods akin to “Linux Ricing” (Cowell, 2016) to content moderation on a decentralized social media platform. While there are many different ways a user can hold or surrender control in the context of today’s digital technologies, we choose to focus on one particular application of subsidiarity in this paper: subsidiarity in terms of data collection. In this way, we evaluate the tradeoffs between privacy and convenience called into question with surveillance capitalism today.
@@
  • RQ1: To what extent is the software we use every day collecting data beyond expected or necessary usage? In other words, what level of control is the software currently operating at?
@@
  • The empirical contributions of our data flow analysis, which provide insight into how various applications (including operating systems and software applications not well-studied in previous literature) could be transmitting our personal data
  • The empirical contributions of our user interviews, which provide insight into what level of data control computer users view as necessary across different types of software applications and why
@@
Beyond furthering the conversation regarding the ethics of data privacy, we hope this paper will demonstrate the principle of subsidiarity as an effective critical lens through which to promote user flourishing in technology design. The principle of subsidiarity can help designers concerned with user well-being promote user autonomy as much as possible, while also being able to discern and justify instances of higher-level control for the benefit of the user.
@@ 2. Related Work
We situate our research within a number of areas of prior work. First, we look at previous work in human-computer interaction that engages with the principle of subsidiarity or similar frameworks. Next, we consider previous research in data collection, sharing, and tracking, especially in the contexts of devices and browsers. Finally, we consider prior research of user perceptions and attitudes towards data privacy. In all cases, we highlight the unique contributions of our work.
@@ 2.1. The Principle of Subsidiarity and Similar Frameworks in HCI and Beyond
Two prior works in HCI have considered the principle of subsidiarity explicitly. Both demonstrated subsidiarity as an effective principle for re-imagining more ethical social media. First, Hasinoff and Schneider show subsidiarity can help improve the balance between context and scale in content moderation (Hasinoff and Schneider, 2022). We build on this work by bringing subsidiarity into a new domain and through conducting empirical studies. In their work on common good-oriented design patterns for social media, Conwill et al. argued that designing for the principle of subsidiarity would encourage more respectful and authentic conversations, and help to resist exploitative mechanisms from tech companies (Conwill et al., 2025). They also note that subsidiarity entails allowing users to customize their experiences more, giving users greater control over the content displayed in their feeds and the amount of notifications they receive. In this work we expand the notion of control into the realm of data privacy.
The principle of subsidiarity has also been proposed as a critical lens for LLM development. Specifically, a recent issue of the Red Hat Research Quarterly proposed that subsidiarity should be a guiding principle for open-source AI development (Craven et al., 2025). Amid debates of how “open” open-source models should be, the authors propose that subsidiarity can help determine how much access different communities have to model components based on how much control of the model they need for their use case (Craven et al., 2025). In this way, subsidiarity is upheld when control is given to the most localized community that is able to effectively make their own decisions.
Related to subsidiarity in HCI is the Anarchist HCI movement (Keyes et al., 2019). Social anarchists hold that “human dignity is greatest when human autonomy is greatest” and that communities are the most appropriate decision-makers and loci of power (Keyes et al., 2019). In this we see resonance with subsidiarity, which also holds that autonomy supports human dignity and that lower levels of decision-making are more just. An even stronger connection between anarchy and subsidiarity lies in how subsidiarity can inform understandings of anarchy. As an example, the Catholic Worker Movement, an anarchist Catholic social justice movement, uses the principle of subsidiarity to define its understanding of anarchy (Linder, 2024). This connection illustrates how our use of subsidiarity as a critical lens for HCI extends previous work in anarchist HCI. Previous contributions within anarchist HCI proposed how anarchism can influence HCI as a discipline, offering ways to re-think the contributions of our works and how HCI researchers relate to their study participants and the broader HCI research community (Keyes et al., 2019). In contrast, our work considers how subsidiarity can influence the design and use of the computer technologies themselves.
@@ 2.2. Data Sharing and Tracking at the Device Level and Browser Level
Prior work in both HCI and other computer science disciplines has investigated privacy and data tracking on different types of devices. These studies have found that various types of devices — fertility trackers (Hudig and Singh, 2025), Apple mobile devices (Bourdoucen and Lindqvist, 2024), and Android devices (Gamba et al., 2020) — access and share data in ways unexpected to the user. This data sharing spans countries and organizations, including to advertising and tracking services (Hudig and Singh, 2025; Gamba et al., 2020). We extend this prior work both by investigating the data flows from different operating systems, specifically Windows and Ubuntu, and analyzing our findings through the critical lens of the principle of subsidiarity. We draw methodologically from the work of Hudig and Singh (Hudig and Singh, 2025), incorporating both data flow monitoring and user interviews in our work.
@@
Google, Facebook, Microsoft, and Twitter are the most dominant third-party trackers, especially in Western countries, with Google being the significantly biggest offender (Schelter and Kunegis, 2021; Gill et al., 2013; Roesner et al., 2012; Lerner et al., 2016; Karaj et al., 2018; Englehardt and Narayanan, 2016; Krishnamurthy and Wills, 2009; Cassel et al., 2022). More recently, TikTok has become a major third-party tracker (Munir et al., 2023). We will rely on these findings to help determine which platforms are “known data harvesters,” which is one piece of our determining the level of control of an application.
Mainstream browsers experience 70% more third-party tracking and advertising requests on average compared to privacy-focused browsers and experience fingerprinting on significantly more websites (Cassel et al., 2022). As such, we include one privacy-focused browser, Brave, in our investigation.
We extend this prior research first by studying Windows and Ubuntu, which do not appear to be studied in the previous literature. Second, a number of the web applications we investigated, especially BlueSky, Discord, and Slack, do not appear in the previous literature.
@@ 3. Theory: The Principle of Subsidiarity
In this section we will further explain the principle of subsidiarity, outline its history and use, and argue why we believe it is an effective critical lens through which to envision a better relationship with technology in the future.
@@
The principle of subsidiarity states that local governing bodies should be empowered to make decisions for themselves, and central authorities should only perform tasks that cannot be done at a local level (for Justice and Peace, 2004). The principle of subsidiarity promotes human dignity through a prioritization of human agency and autonomy: by encouraging decisions to be made at the lowest level possible, individuals and communities are able to make decisions about matters that affect them most, fostering a sense of responsibility and participation (Benedict XVI, 2009). However, subsidiarity is not libertarianism: an important element of subsidiarity is acknowledging the importance of higher levels of intervention when necessary and with appropriate justification (for Justice and Peace, 2004).
@@
Historically, the principle of subsidiarity is perhaps most prominently associated with the Christian Democratic political ideology, having its roots in both Catholic Social Teaching (for Justice and Peace, 2004) and the neo-Calvinist teaching of sphere sovereignty (43). However, its resonance goes beyond Christian Democrats alone. For example, today subsidiarity may be most well-known as a general principle of European Union law (43). Importantly, the principle of subsidiarity can resonate with both the political left and the political right: this broad resonance makes it worthy of our consideration. The political left can find resonance with the principle of subsidiarity because it upholds participatory government, resists centralized economic and political power, and can support self-determination for historically marginalized groups. Subsidiarity can resonate with the political right because it can justify limited government and supports giving primacy to traditional institutions like the family in addressing moral questions.
@@
While typically associated with questions of political organization, subsidiarity can apply in other contexts. For example, in the workplace, a boss can empower her employees through the principle of subsidiarity by avoiding micro-managing. In this way, the gifts and talents of the employees are honored by giving the employee a greater ability to contribute freely.
@@
Subsidiarity can offer much in the space of ethical technology design as well. Importantly, Catholic Social Teaching — the Catholic Church’s doctrine about social justice and the primary framework through which subsidiarity emerged — began in response to the societal concerns of the Industrial Revolution (for Justice and Peace, 2004). Specifically, subsidiarity emerged as a response to increasing centralization of economic power (for Justice and Peace, 2004), a trend we see repeating today with “big tech.” While systems like surveillance capitalism undermine personal autonomy, subsidiarity aims to reclaim it. Yet, because subsidiarity guides the discernment of when intervention from higher levels of power is necessary, it can help to tease out the tension between user desires for privacy and the convenience of giving up their data, and help to clarify when data collection is appropriate or necessary.
@@
As mentioned in the introduction, subsidiarity in a technical context means control of digital spaces should be made at the lowest level possible and the highest level necessary. In the specific application to data collection, subsidiarity means that data collection should occur at the lowest level possible and highest level necessary for effective use of the application. We identify the levels of control we consider in this paper from lowest to highest as follows: user, community, and platform. Control at the user level means that data collection should be transparent and require informed consent to be collected and used in ways not immediately obvious from the primary function of the application. Control at the platform level means that companies can collect our data how they please, typically in exchange for the conveniences afforded by them using and/or sharing our data. Finally, control at the community level pertains to communities, employers, organizations, or special interest groups having control of their participants’ data, typically for the greater benefit of that community. This may include an employer providing software to its employees, a high school providing Chromebooks to its students through the Google for Education program, or a community-run online forum for a special interest or hobby, among other examples.
@@ 4. Methodology
Prior work has highlighted that security and privacy concerns vary across different stages of the data life cycle: data collection, usage, and retention (Ali et al., 2025). In our methodology we focus on subsidiarity in relation to data collection. We focus on data collection because it is both the gateway to the other stages and also where we find the user has the most potential for control in the current technical landscape: a user can, to some extent, control if their data is collected through choosing to use or reject particular platforms that are known to collect user data. However, once data is collected it is less straightforward for the user (at least in the current technical landscape) to control how their data will be used or retained by the platform. It is also the most straightforward stage for us to study: we can monitor if platforms are collecting data but once data is collected we cannot know for sure how they are using it. Additionally, even if platforms provide mechanisms for users to “delete” their data, in some cases this data may actually be retained by the platform (Ng, 2019), making the study of data retention more difficult.
@@ 4.1. Data Flow Monitoring
We monitored the network transmissions of different types of everyday software platforms in order to infer the extent to which these platforms are collecting and sharing data. Although it is impossible to get a full picture of what is going on, our aim was to get as much of a sense as we could of the entities being contacted and the potential endpoints where user data could be being transmitted. To do so, we used the network analysis tool Wireshark (3) to capture the network activity while using different software platforms. We investigated the network activity of multiple operating systems and browsers on their own, as well as various web applications, run in a particular browser on a particular operating system.
@@ 4.1.1. Operating System and Browser Selection
We wanted to compare operating systems and browsers that are more conventional — i.e., owned by large tech companies, who do not make many promises about privacy — to operating systems and browsers that are open-source and/or claim to better protect user privacy. Thus, we selected the operating systems Microsoft Windows and Ubuntu Linux, and the browsers Google Chrome, Firefox, and Brave (1). Brave has been demonstrated in the literature to be the most effective browser at reducing third-party tracking and advertising requests (Cassel et al., 2022).
@@ 4.1.3. Data Collection
The social networking experiments ran as follows: open the browser, navigate to the site, log in, scroll for one minute, log out, and close the browser. The protocol was slightly different for YouTube: for that experiment we did not log in, we searched for “Taylor Swift,” clicked on the first video, and watched for one minute.
Finally, the chat experiments ran as follows: open the browser, navigate to the site, log in, send 3-5 short (between 2 words and one sentence long) messages to the same recipient, and close the browser. (The recipient did not respond during the course of the experiment.)
@@ 4.1.4. Data Analysis
Through data flow monitoring, because we are observing from the outside, we cannot know if data is only sent to the application for the purpose the user intended (e.g., sending email data to a mail application for the purpose of sending an email, no other data is collected from the user, and the data from the email itself is not used for any other purposes), or if the data sent for an intended purpose is additionally being collected for purposes unintended by the user. Additionally, it is harder to distinguish between community and platform levels of control in data flow monitoring, because that is determined more by who the data is being shared with rather than that data is shared at all. Thus, we cannot definitively determine if the software operates at user, community, or platform levels of control. Instead, we determine if the software operates at a lower (closer to user-level) or higher (closer to platform or community-level) of control through the following criterion: is application data being sent to the parent company (or any other company) while the software platform is running idle? Additionally, are the platforms only contacting IP addresses affiliated with their parent company and/or known cloud service providers, or are they contacting unexpected IPs (e.g., advertisers or trackers)? Finally, is the parent company a known data harvester (e.g., Google, Meta)? We define the criteria for a “known data harvester” as any company for which it is widely known (Schelter and Kunegis, 2021; Gill et al., 2013; Roesner et al., 2012; Lerner et al., 2016; Karaj et al., 2018; Englehardt and Narayanan, 2016; Krishnamurthy and Wills, 2009; Cassel et al., 2022; Munir et al., 2023) that their business model relies on the harvesting of user data to sell to third parties or to share with third-party advertisers. If the software platforms we examine are doing any of these things we determine that they are operating at a higher level of control; if none, they operate at a lower level of control. The flow chart in Figure 1 outlines our decision-making path for the level of control a software application operates at.
@@
We determined both the IPs being contacted during a particular experiment and the amount of application data being sent to various IPs through the Wireshark logs. We extracted the IP addresses from the Wireshark logs and used the reverse IP address lookup tool IPAPI (24) to determine the organizations associated with those IP addresses. We focused on the destination IP addresses, that is, the IPs being contacted by the software. We were unable to read the contents of the packets due to encryption, which is a common challenge in work on network monitoring (Hudig and Singh, 2025).
@@
Finally, for the activity experiments, because we would expect application data to be transferred to the parent company while the user is performing an activity, we primarily focus on if the IPs contacted are expected or unexpected.